### v1

# PsyMem: Fine-grained psychological alignment and Explicit Memory Control for Advanced Role-Playing LLMs

Abstract
Existing LLM-based role-playing methods often rely on superficial textual descriptions or simplistic metrics, inadequately modeling both intrinsic and extrinsic character dimensions. Additionally, they typically simulate character memory with implicit model knowledge or basic retrieval augment generation without explicit memory alignment, compromising memory consistency. The two issues weaken reliability of role-playing LLMs in several applications, such as trustworthy social simulation. To address these limitations, we propose PsyMem, a novel framework integrating fine-grained psychological attributes and explicit memory control for role-playing. PsyMem supplements textual descriptions with 26 psychological indicators to detailed model character. Additionally, PsyMem implements memory alignment training, explicitly trains the model to align character’s response with memory, thereby enabling dynamic memory-controlled responding during inference. By training Qwen2.5-7B-Instruct on our specially designed dataset (including 5,414 characters and 38,962 dialogues extracted from novels), the resulting model, termed as PsyMem-Qwen, outperforms baseline models in role-playing, achieving the best performance in human-likeness and character fidelity.
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
1 Introduction
Recent advancements in large language models (LLMs) (Achiam et al., 2023; Anthropic, 2024a; Anil et al., 2023) have unlocked new possibilities for implementing role-playing systems through language-based behavior simulation (Shanahan et al., 2023). Role-playing systems demonstrate societal value across applications like interactive gaming (Wang et al., 2023a) and AI counseling (Zhang et al., 2024), particularly in enabling social simulation. Such simulation allows modeling of human interactions and societal dynamics (Park et al., 2023), offering insights into policy design and conflict mediation (Zhou* et al., 2024). To achieve trustworthy social simulation, LLM-based role-playing must preserve persistent traits, memories, and behavioral logic to ensure agents reliably mirror specified attributes in evolving social contexts.
Recent works in social or group behavior simulation (Kosinski, 2024; Park et al., 2023; Zhou* et al., 2024) often rely on general LLMs for role-playing, using context-based learning (Mann et al., 2020) and instruction follow-up (Ouyang et al., 2022). However, these approaches (Yu et al., 2022) commonly struggle to precisely control character attributes due to the unsatisfactory role-playing ability of general LLMs, especially for ordinary characters, as Fig.1 shows. Some recent works (Shao et al., 2023; Zhou et al., 2023; Tao et al., 2023) proposed the LLMs that are dedicated to role-playing to improve character attribute consistency. However, there are still the following two issues: 1) Oversimplified characterization and 2) Weak memory control.
The oversimplified characterization issue. From both neuroscientific and psychological perspectives, human attributes can be described through two broad dimensions: intrinsic attributes and extrinsic attributes (Snyder, 1983; Kahneman, 2011). Intrinsic attributes, such as personality traits and values (Ravlin and Meglino, 1987; Stumpf and Dunbar, 1991), play a significant role in decision-making, particularly in long-term decisions (Borghans et al., 2008; Roberts et al., 2008). On the other hand, extrinsic attributes (Kahneman, 2011), such as behavior, language, and body language, reflect an individual’s intuitive responses in immediate situations (Bargh et al., 1996).
However, existing role-playing LLMs typically rely on basic textual descriptions or narrow metrics to define the target role, which fails to capture the full complexity and depth of the character. For example, the work (Zhou et al., 2023) defines character attributes using textual descriptions based on seven dimensions (including identity, interests, experiences, and other relevant traits). Some recent research (Ran et al., 2024; Yu et al., 2024; Cui et al., 2023) has attempted to use MBTI for personality trait evaluation to improve character control. However, MBTI employs a binary approach to assess personality traits, which has low test-retest reliability (Stein and Swan, 2019; Hunsley et al., 2015) and cannot precisely quantify attributes (Nowack, 1996), such as how introverted a person is.
Figure 1: The performance comparison is conducted on two subsets: “Ordinary,” consisting of 20 randomly selected characters from our test set, and “Famous,” consisting of 20 well-known characters. The average attribute scores are calculated as the mean of the quantized scores across the 20 characters in each subset.
The weak memory control issue. Numerous neuroscientific studies have shown that(Stumpf and Dunbar, 1991; Ravlin and Meglino, 1987; Pickering and Garrod, 2004), during decision-making, the brain relies on not only personality traits (such as emotional responses and values) but also long-term memory, particularly in fast-paced decision-making contexts. However, existing role-playing models (Shao et al., 2023; Zhou et al., 2023) typically rely on the model’s internal knowledge to simulate character memory, which works for famous characters but often fails to provide memory for less familiar or unseen ones. Retrieval-Augmented Generation (RAG) (Ram et al., 2023) has been proposed as a promising solution (Zhong et al., 2024), but we observe that simply integrating RAG within Role-Play LLMs’ inference stage does not adequately address memory management in the context of role-play (see Table 2).
In this work, we propose a widely recognized modern psychological evaluation framework (Costa and McCrae, 2008; Schwartz, 1992; Thomas, 2008; Zimbardo and Ruch, 1975) and a graph-structured memory mechanism to construct the PsyMem framework, enabling LLMs to achieve exceptional role-playing abilities. Drawing from modern psychology and neuroscience, we enhance the description of character attributes by leveraging both latent psychological attributes and explicit behavioral patterns (Snyder, 1983; Kahneman, 2011), with personality and values corresponding to latent psychological attributes, and behavioral decision-making, social interaction, and leadership representing explicit behavioral patterns. Overall, we describe characters using 26 quantitative indicators and a small amount of textual descriptions. We extract real data from novels and construct a publicly available role-playing dataset. As shown in Table 1, this dataset contains 5,414 established characters, 38,962 conversations, and 536,636 utterances, much larger than existing role-playing datasets.
For character memory, inspired by the hippocampus and cognitive mapping mechanisms (Eichenbaum, 2015; Hartley et al., 2014; Tacikowski et al., 2024; Tavares et al., 2015), we transform characters, events, and relationships into a knowledge graph. Unlike previous role playing work (Edge et al., 2024) directly using implicit model knowledge or integrating role memory in inference, we incorporate memory information into the training process, forcing the model to learn how to respond given retrieved memory. This approach effectively enhances the model’s memory consistency, ensuring that role-playing model generates responses based on not only character attributes but also memory.
In general, our five main contributions are as follows:
  • We propose PsyMem, a novel LLM-based role-playing framework that establishes precise control over character behaviors through the synergistic integration of fine-grained psychological profiling (e.g., personality, value) and dynamic memory schemata.
  • Beyond textual descriptions, we present a comprehensive set of attributes that bridges intrinsic (e.g., values, emotions) and extrinsic (e.g., behaviors, social interactions) character dimensions, operationalized via 26 measurable indicators. This attribute set captures both the internal motivations and external behavioral manifestations of characters.
  • We introduce a memory alignment training paradigm that establishes explicit cognitive anchoring between role-playing responses and character memory. This methodology enhances role-playing through a dual-control mechanism, enabling precise regulation of generative outputs via both trait-based attributes (e.g., personality dimensions) and episodic memory constraints.
  • We introduce the use of role-based general supervised fine-tuning data (transforming general SFT data into role-playing style SFT data), demonstrating that incorporating such data effectively enhances the performance of Role-Playing LLMs.
  • Extensive experiments demonstrate that the proposed PsyMem framework significantly improves the role-playing ability of LLMs. For instance, by training Qwen2.5-7B-Instruct on our dataset (introduced in Section 3) specially designed for PsyMem, the resulting model, termed as PsyMem-Qwen, outperforms baseline models in role-playing, achieving the best performance in human-likeness and character fidelity.
2 Related Work
2.1 LLMs-based Role-Playing
In existing role-playing works, the main approaches are Nonparametric Prompting and Parametric Training (Chen et al., 2024). In Nonparametric Prompting, Yu et al. (2022) use carefully designed character descriptions to control general LLMs for role-playing without fine-tuning. However, they face significant challenges in accurately reflecting the intrinsic relationships between character profiles and dialogue content.
In Parametric Training, Chen et al. (2023) focus on mimicking characters from the Harry Potter series, while Wang et al. (2023b) introduced the first role-playing dataset with 100 characters generated using GPT-3.5 prompts. Subsequent works (Li et al., 2023; Zhou et al., 2023; Shao et al., 2023) have expanded and enriched these datasets by improving the number of characters and the diversity of dialogue scenarios, all leveraging various GPT models. Meanwhile, Lu et al. (2024) adopted a self-alignment approach to dataset construction, moving away from the previous method of cheaply imitating the role-playing capabilities of GPT models. Zhou et al. (2023) leveraged GPT-4 to extract character dialogues from a diverse collection of Chinese novels and scripts. More recently, Yu et al. (2024) focused on addressing the bias between predefined characters and specific scenario dialogues, and constructed a dataset of 311 characters.
Our work belongs to parametric Training. We argue that a character’s behavior is jointly influenced by their attributes, memory, and the scenario. Drawing from modern psychology, we segment character attributes using contemporary psychological frameworks and construct a dataset containing 38,962 conversations (536,636 utterances) from real novels, leveraging both character attributes and memory to enhance the model’s role-playing ability.
2.2 Memory Mechanism in Role-Playing.
Existing LLMs-based Role-Playing commonly relies on the internal knowledge within the model weights to implicitly mimic character memory (Zhou et al., 2023; Lu et al., 2024; Li et al., 2023). For example, Shao et al. (2023) and others push the boundaries of knowledge to protect character memory, while Lu et al. (2024) and others use a self-alignment approach to distinguish the different memory backgrounds of various characters. Relying on the internal knowledge within the model is effective for famous characters but fails for infamous and unseen characters. Some works apply Retrieval-Augmented Generation (RAG) (Ram et al., 2023) which dynamically retrieving external contexts to explicitly mimic character’s memory. For example, Zhong et al. (2024) adopts a memory update mechanism inspired by Ebbinghaus’s forgetting curve theory, emphasizing the recall of relevant memories to understand the user’s personality. However, as shown in Table 2, existing role-playing models tend to perform better when simulating famous characters, while their performance significantly drops for ordinary characters that rely on external memory.
Unlike previous works, we integrate explicit memory control into the training framework of role-playing models. This approach encourages the model to strictly rely on actual retrieved memories during dialogue generation in inference.
2.3 Role-playing Evaluation.
The current evaluation process consists of dialogue generation and dialogue assessment. Due to its efficiency and completeness, automated dialogue generation is widely adopted. Evaluation methods are generally categorized into three types: Metric-based Evaluation(Li et al., 2023; Wang et al., 2023b), Human Evaluation(Zhou et al., 2023), and “LLMs as Judges”(Shao et al., 2023; Lu et al., 2024). Metric-based evaluation primarily measures the model’s alignment with standard responses, such as original text from novels or manually annotated content. While human evaluation is more precise, its high cost and lack of reproducibility limit its scalability. In contrast, “LLMs as Judges”, with its efficiency, low cost, and strong scalability, is increasingly becoming the preferred choice.
3 Dataset Architecture
An increasing number of role-playing datasets (Ran et al., 2024; Cui et al., 2023) have begun incorporating modern psychological indicators into their characterization frameworks. However, previous studies typically introduced only a limited selection of psychological attributes as supplementary descriptors, without comprehensively integrating psychological systems into character portrayals (Yu et al., 2024; Wang et al., 2024). Moreover, existing datasets predominantly focus on character attributes, overlooking the significant influence memory has on characters’ behaviors, thought processes, and speech styles(Stumpf and Dunbar, 1991; Ravlin and Meglino, 1987; Pickering and Garrod, 2004). This limitation becomes particularly evident when role-playing characters possess limited background knowledge during the pretraining stage (see Figure 1).
To address these shortcomings, we design a novel dataset with the architecture formulated as
[…]
where $N$ is the number of data sample. Each sample is a dialogue turn where $R_{i}$ is the role for response; $P_{i}$ is the profile of $R_{i}$, featuring 26 quantitative psychological dimensions; $C_{i}$ is the previous dialogue turns before the current dialogue. $x_{i}$ is the query to $R_{i}$ and $y_{i}$ is the corresponding response. $M_{i}$ is the retrieved memory response of role $R_{i}$ with respect to query $x_{i}$ and the previous dialogue $C_{i}$.
In this work, the dataset $D_{\mathrm{RP}}$ contains 38,962 samples, extracted from 539 novels characterized by a wide range of genres, rich content, and narrative perspectives of third person.(The genre distribution is presented in Figure 2.)
Dataset Character Conversation Profile Style
Num. SA Conv. Avg. Turns Auth. PsyFrm QI Text Desc. Expl. Mem.
HPD 113 1,191 13.2
CharacterGLM 250 1,034 15.8
RoleLLM 100 140,726 2
CharacterLLM 9 14,300 13.2
ChatHaruhi 32 54,726 >2
DITTO 4002 7,186 5.1
Beyond Dialogue 311 3,552 6.5
CoSER 17,966 29,798 13.2
MMRole 85 14,346 4.2
CharacterBench 3956 13,162 11.3
CharacterEval 77 1,785 9.3
Ours 5,414 38,962 13.8
Table 1: Dataset Statistics. Comparison between our dataset and existing open-source role-play datasets. Columns are categorized into: Character (Num.: number of character roles; SA: scene alignment), Conversation (Conv.: number of dialogues; Avg. Turns: average dialogue length; Auth.: whether dialogues are fully sourced from real, non-generated conversations), and Profile Style (PsyFrm: adherence to a psychological framework; QI: presence of quantitative indicators; Text Desc.: availability of textual character descriptions; Expl. Mem.: support for explicit memory retrieval during interaction).
Figure 2: The genre distribution in the dataset.
Figure 3: (a): The fine-grained character attributes designed in this work. (b): The two stages training of role-playing LLM. First, the model learns basic role-playing without character memory. In the second stage, we dynamically retrieves memory relevant to the current query and dialogue context from graph-structured character memory. We then integrate character profile, retrieved memories, dialogue history, and current input to enhance role-playing precision, training the model to align responses with both fine-grained character profile and contextual memory. (c):We assess the role-playing LLM by first generating a multi-turn dialogue (up to 15 turns) between two designated roles, followed by three rounds of scoring with GPT-4o based on a quantitative rubric, and report the mean score.
3.1 Quantifiable Character Attributes
We systematically construct quantifiable character attribute dimensions aligned with contemporary psychological frameworks by categorizing character traits into latent psychological attributes and explicit behavioral patterns (Kahneman, 2011; Snyder, 1983), as illustrated in Figure 3a. These two dimensions distinctly influence individual behavior, communication styles, decision-making processes, and interpersonal interactions. Besides, a small amount of text describing the social attributes of the characters is also included as a supplement.
Latent psychological attributes reflect internal, often subconscious aspects of character, guiding motivations, preferences, and emotional responses. To describe these latent psychological attributes comprehensively, we utilize the following psychological frameworks:
1) Personality: Evaluated using the Big Five Personality Model (Costa and McCrae, 2008), encompassing stable emotional, cognitive, and behavioral tendencies. Traits such as openness, conscientiousness, extraversion, agreeableness, and neuroticism directly influence communication styles, social interactions, and general life choices.
2) Values: Assessed through Schwartz’s Theory of Basic Values (Schwartz, 1992), which classifies core beliefs into ten universally recognized types. These values deeply affect decision-making, priorities, ethical judgments, and interpersonal alignment, enriching the psychological depth and motivational dimensions of individuals.
Explicit behavioral patterns manifest externally in observable actions, interactions, and communication styles. We capture these explicit patterns using:
1) Social & Leadership Styles: Measured through six dimensions based on Zimbardo’s psychological principles (Zimbardo and Ruch, 1975), significantly influencing observable interpersonal behaviors, communication effectiveness, leadership approaches, and social dynamics in groups.
2) Behavioral Decision: Analyzed using the Thomas-Kilmann Conflict Mode Instrument (TKI) (Thomas, 2008), emphasizing decision-making processes when characters face conflicts or challenges. This explicitly captures how emotional, social, and heuristic influences shape specific decisions, providing precise and actionable insights into behavioral responses.
3.2 Character Memory Integration
Neuroscientific research underscores the significant role memory plays in shaping human behavior, communication, and decision-making processes (Stumpf and Dunbar, 1991; Ravlin and Meglino, 1987; Pickering and Garrod, 2004). Memories influence not only how individuals perceive and interpret events but also guide their interactions and responses within social contexts, highlighting the importance of accurately capturing and utilizing memory in character portrayal.
Therefore, we introduce character memory data $M$ to the role playing dataset to enable the memory simulation capability of role playing model. Specifically, we first build a knowledge graph based on the novel via GraphRAG (Edge et al., 2024), and then retrieve relevant memory for the current dialogue via entity mentions, relational triples, and raw dialogue chunks associated with specific scenes. Memory alignment is evaluated on two key dimensions: correctness and rationality. Correctness assesses whether the model accurately utilizes stored character information in response to queries, while rationality examines logical coherence and contextual appropriateness. Additionally, simple summaries or vague memory references are identified as non-rational uses of memory. To further enhance robustness, irrelevant memory items are deliberately introduced as noise, promoting stable, realistic, and consistent character portrayals reflective of human memory dynamics.
In addition, we define character-specific knowledge boundaries to ensure consistency—for example, Quasimodo from The Hunchback of Notre-Dame would not be aware of torch. Based on the absence of relevant retrievable memories, we used GPT-4o to generate 800 refusal-style QA pairs for 100 characters, grounded in real-world constraints and narrative context. These were integrated into novel-based dialogues to maintain coherence and character fidelity.
3.3 Role-based SFT Data
We propose a Role-based general supervised fine-tuning (SFT) data construction approach. Dialogue samples from the Pure-Dove (PD) dataset (Daniele and Suphavadeeprasit, 2023) are grouped randomly, and LLMs infer each speaker’s latent psychological attributes and explicit behavioral patterns. Scene-specific information is dynamically generated using fine-tuned LLMs based on dialogue contexts, significantly improving character consistency and alignment during interactions, as demonstrated by our experimental evaluations.
Similar to Equation 1, we construct a dataset defined as:
[…]
Where $R_{i}$ denotes the role responsible for the response, and $P_{i}$ is the profile of $R_{i}$ defined by 26 quantitative psychological dimensions. $C_{i}$ represents the preceding dialogue context, $x_{i}$ is the current query directed to $R_{i}$, and $y_{i}$ is the corresponding response. These $(x_{i},C_{i},y_{i})$ triples are drawn from general supervised fine-tuning datasets. In addition, $M_{i}$ refers to the retrieved memory response relevant to role $R_{i}$ with respect to $x_{i}$ and $C_{i}$.
4 Role Playing Training and Evaluation
4.1 Training
In this work, we transfer general LLMs to role-playing models via supervised fine-tuning. Role-playing presents a unique challenge for language models: maintaining both general capabilities and character-specific attributes simultaneously. This challenge exemplifies the fundamental tension in transfer learning between preserving general knowledge and specializing for domain-specific tasks. Traditional fine-tuning approaches (Kotha et al., 2024) often suffer from catastrophic forgetting, where optimization for character fidelity compromises the model’s general capabilities. Inspired by the Dual-stage Mixed Fine-tuning (DMT) (Dong et al., 2023) method, we introduce a two-stage training strategy for role-playing, effectively balance character fidelity and the preservation of general language understanding capabilities, as illustrated in Figure 3b.
Specifically, we introduce a two-stage progressive training framework to systematically develop LLMs into specialized Role-Playing Models: Stage 1 establishes core role-playing capabilities through fine-tuning on data emphasizing fundamental role characteristics. Stage 2 enhances specialization via memory-augmented fine-tuning, incorporating long-term memory contexts and general role-based information. This approach ensures a balanced development of character-specific attributes while simultaneously preserving the model’s general capabilities.
Stage 1: Foundational Role-Playing Capacity Development.
In this stage, we train LLMs to learn foundational role-playing ability based only on quantifiable attributes and basic information. We fine-tune the model using a subset $D_{\mathrm{RP1}}$ sampled from the overall role-playing dataset to establish its fundamental role-playing abilities. The objective function is formally defined as:
[…]
where ()($\cdot$)( ⋅ ) in $D_{\mathrm{RP}}$ represents the tuple $(R,P,C,x,y)$, $D_{\mathrm{RP1}}$ is a subset sampled from the overall role-playing dataset $D_{\mathrm{RP}}$, excluding the retrieved memory $M$.
Stage 2: Integrated Memory-Augmented Specialization. To achieve integrated memory-augmented specialization, we further specialize the model by fine-tuning it on a distinct subset from the role-playing dataset $D_{\mathrm{RP}}$, augmented with long-term memory contexts retrieved via graph-based methods, and combined with general role-based SFT data $D_{\mathrm{RB}}$. This stage employs a weighted combination of two loss terms, explicitly designed to mitigate forgetting of general capabilities while simultaneously enhancing the model’s ability to effectively leverage long-term memory contexts:
[…]
where $\alpha=20$.
The two loss terms are individually defined as follows:
[…]
where ()($\cdot$)( ⋅ ) in $D_{\mathrm{RP2}}$ represents the tuple $(R,P,M,C,x,y)$ and in $D_{\mathrm{RB}}$ represents $(R,P,C,x,y)$. $M$ denotes the retrieved memory response for role $R$ regarding the query $x$ and prior dialogue context $C$, obtained via graph-based retrieval methods and subsequently filtered by LLMs.
Model Character Fidelity Character-independent
Per. Val. SL BD Mem. Avg. H-like Cons. Coh.
General LLMs
GPT-4o 79.08 80.11 77.43 73.29 94.40 80.86 66.60 89.28 99.20
GPT-3.5-Turbo 79.08 79.42 74.15 72.11 84.60 77.87 34.40 88.60 96.20
Claude-3.5-sonnet 80.34 80.29 78.51 74.57 91.80 81.10 75.00 89.50 99.20
Qwen-Max 78.67 80.42 76.54 72.37 90.60 79.72 57.60 89.48 97.60
Yi-Large 80.20 80.09 78.27 76.82 92.00 81.48 83.80 88.18 98.40
Deepseek-R1 80.98 81.38 77.88 77.12 95.00 82.47 77.20 87.57 96.20
Deepseek-V3 79.82 80.33 79.16 75.08 93.40 81.56 76.40 89.04 99.00
Llama3.1-8B-instruct 79.17 80.14 76.76 76.01 83.20 79.06 64.20 88.74 99.00
Qwen2.5-7B-Instruct 77.44 80.02 76.15 76.51 85.60 79.14 64.40 89.58 97.40
Role-play LLMs
CharacterGLM-6B 80.66 80.67 79.35 71.20 40.20 70.42 52.80 89.61 73.00
Baichuan-NPC-Turbo 76.53 79.05 74.29 72.27 87.20 77.87 61.60 90.36 96.40
Hunyuan-Role 79.31 79.71 78.24 78.38 82.00 79.53 87.40 89.16 91.20
Higgs-Llama3-70B 80.02 80.61 75.77 74.08 93.20 80.74 60.40 88.40 99.60
CoSER-70B 81.21 81.90 78.83 76.35 93.00 82.28 81.80 89.44 97.20
PsyMem-LLama (ours) 80.47+1.30 81.93+1.79 78.57+1.81 74.47-1.54 90.20+7.00 81.13+2.07 85.20+21.00 89.93+1.19 95.60-3.40
PsyMem-Qwen (ours) 80.40+2.96 81.74+1.72 80.80+4.65 78.48+1.97 91.80+6.20 82.64+3.50 87.60+23.20 89.64+0.06 96.20-1.20
Table 2: Model evaluation across multiple criteria. Bold indicates best performance. Superscripts show change from base model. Abbreviations: Per. = Personality, Val. = Values, SL = Social & Leadership, BD = Behavioral Decision, Mem. = Memory, H-like = Human-likeness, Cons. = Consistency, Coh. = Coherence
4.2 Evaluation
We evaluate role-playing capabilities of LLMs by adopting the widely used “LLMs as Judges” approach (Shao et al., 2023; Lu et al., 2024), as illustrated in Figure 3c. The evaluation process is structured into two main categories: Character-independent Capabilities and Character Fidelity. Additionally, in the Tabel 4, we further discuss the discrepancies between LLM-based judgments and human evaluations across 26 quantitative indicators to validate the effectiveness of the “LLMs as Judges” approach.
4.2.1 Character-independent Capabilities
Following previous research(Chen et al., 2024; Yu et al., 2024), we measure Character-independent Capabilities using three metrics: Human-likeness, Consistency and Coherence. In the Human-likeness task, we use GPT-4o with few-shot examples to determine whether a dialogue was generated by a human, ultimately computing the accuracy score. For consistency, we evaluate the coherence between different parts of the conversation to assess whether the role-playing remains consistent throughout long dialogues. Additionally, for coherence, we assess how logically connected and contextually appropriate the responses are across the entire dialogue, ensuring the conversation flows naturally.
4.2.2 Character Fidelity
Character Fidelity assesses how accurately an LLM portrays specific role-playing characters. As described in Section 3.1, we employ modern psychological quantification frameworks defined in our dataset design, covering four primary dimensions: Personality (Big Five Personality Model), Values (Schwartz’s Theory of Basic Values), Social and Leadership Styles (Zimbardo’s psychological principles), and Behavioral Decision-making (Thomas-Kilmann Conflict Mode Instrument, TKI). These dimensions are quantified into numeric scales using standardized psychological evaluation methods. Additionally, we separately evaluate the model’s memory, focusing on the correctness and rationality of memory alignment. The Memory dimension is measured using accuracy scores derived from a true/false evaluation format.
To calculate similarity between role-playing characters, we employ Normalized Euclidean Distance (NED), computed as follows:
[…]
where $\Delta X=X_{\max}-X_{\min}$, $n$ is the number of subdimensions, $X_{\mathrm{ref}}$ is the standardized reference value, and $X_{i}$ represents the measurement for the $i$-th subdimension, with possible values from $Z=\{1,2,3,4,5\}$. This formulation is robust, as it normalizes the mean squared deviation across different scales, yielding a more interpretable similarity metric.
5 Experiments
In this section, we explain the dataset, baselines, and metrics design in the experimental setup, and discuss the main results. Additional ablation experiments are discussed in Section 5.3.
5.1 Experimental Setup
Evaluation Dataset.
To mitigate potential knowledge bias introduced during the pretraining phase, we construct an evaluation dataset for our main experiments by selecting characters and scenes from novels published after June 2024. This design choice aims to ensure that models have had minimal or no exposure to the selected texts during pretraining, thereby reducing reliance on memorized character-specific knowledge. We first score 100 contemporary novels using large language models (LLMs) and then select 50 representative characters spanning diverse genres. For each character, we collect ten scenes directly sourced from the original text.
Baselines.
In our benchmark evaluation, we tested both open-source and proprietary advanced chatbots, each representing the state-of-the-art within their respective frameworks. For general-purpose models, we compared the proposed PsyMem with several prominent baselines, including GPT-4o(Hurst et al., 2024), GPT-3.5-turbo (OpenAI, 2023),Claude-3.5-sonnet(Anthropic, 2024b), Qwen-Max (Team, 2024), Yi-Large (Young et al., 2024), Deepseek-R1(Guo et al., 2025), DeepSeek-V3 (Liu et al., 2024), Llama3.1-8B-instruct(Grattafiori et al., 2024) and Qwen2.5-7B-Instruct (Yang et al., 2024). These models are considered some of the best in the field, offering a broad range of capabilities across various domains. In the category of Role-Playing-focused baselines, we made comparisons with Hunyuan-Role(Tencent Cloud, 2023), Baichuan-NPC-Turbo(Baichuan AI, 2023), and CharacterGLM-6B (Zhou et al., 2023),Higgs-Llama3-70B(AI, 2024),CoSER-70B(Wang et al., 2025), which are specifically designed to handle character-driven interactions and story-based tasks.
Implementations.
With the proposed PsyMem framework, we train our models based on Llama3.1-8B-Instruct and Qwen2.5-7B-Instruct, and denote the trained models as PsyMem-LLama and PsyMem-Qwen, respectively. All models were trained on four A800 80G GPUs for 3 epochs with a global batch size of 128 and a sequence length of 8,192. LoRA was used with a per-device batch size of 2 and 16 gradient accumulation steps. We adopted a cosine learning rate schedule with a warmup ratio of 0.1, and set the learning rate to 1e-4.
Metrics
We utilize GPT-4o to evaluate 50 established characters across 500 scenarios, each scenario comprising 15 dialogue turns. Evaluations are conducted based on the dimensions described in Section 3, with mean scores calculated accordingly.
5.2 Main Results
We present the main results in Table 2. We extracted 500 real novel scenes to evaluate all models, assessing them from two perspectives: Character Fidelity and Character-independent Capabilities. Notably, in terms of Memory alignment, Proprietary models demonstrate a more reasonable approach—not merely summarizing but leveraging memory to portray characters more authentically. Although Higgs-LLaMA3-70B and CoSER-70B exhibit relatively stronger performance in memory handling compared to other role-play-specialized models, there remains a noticeable gap between them and top-performing general-purpose models. Within the General Baselines, we observe that Proprietary models still significantly outperform open-source models. GPT-4o and DeepSeek-V3, as outstanding proprietary models, exhibit impressive performance in Character-independent Capabilities. Additionally, Baichuan-NPC-Turbo, as a role-play expertise baseline model, surpasses general baselines in self-consistency, indicating that this metric is particularly important for role-playing agents. Furthermore, it is worth mentioning that DeepSeek-R1 and CoSER-70B excel in Character Fidelity.
Empirical results under the PsyMem framework reveal that, relative to their original counterparts, PsyMem-Qwen and PsyMem-LLaMA achieve character fidelity improvements of 2.07% and 3.50%, respectively. In addition, PsyMem-Qwen demonstrates strong performance in the Social & Leadership and Behavioral Decision dimensions, achieving scores of 80.80% and 78.48%, respectively. PsyMem-LLaMA, on the other hand, obtains a score of 81.93% in the Value dimension, surpassing all other evaluated models in this category. Moreover, PsyMem-Qwen, despite having only 7B parameters, surpassed all measured baselines in Character Fidelity. Additionally, it is worth noting that both models exhibited remarkable improvements in Human-likeness. Overall, the PsyMem framework has demonstrated strong effectiveness in our benchmark tests, enabling precise character control through quantitative character attribute metrics and memory mechanisms.
Model Character Fidelity
Per. Val. SL* BD* Mem. Avg.
Llama3.1-8B-Instruct
+ PsyMem (w/o Mem.) 80.11 81.90 79.23 76.36 44.20 72.36
+ PsyMem 80.47 81.93 78.57 74.47 90.20 81.13
Qwen2.5-7B-Instruct
+ PsyMem (w/o Mem.) 81.02 81.33 82.10 79.02 53.00 75.29
+ PsyMem 80.40 81.74 80.80 78.48 91.80 82.64
Table 3: Ablation study on memory integration in the PsyMem framework. Per.: Personality, Val.: Values, SL: Social & Leadership, BD: Behavioral Decision, Mem.: Memory.
Metrics Per. Val. SL* BD*
Cosine 0.98 0.96 0.97 0.97
Table 4: Cosine similarity between human and GPT-4o scores. Per.: Personality, Val.: Values, SL: Social & Leadership, BD: Behavioral Decision.
5.3 Ablation Study
Model Character Human Consis- Coher-
Fidelity -likeness tency ence
Llama3.1-8B-Instruct 79.06 64.20 88.74 99.00
     + PsyMem (Original PD data) 80.51 70.80 89.49 82.20
     + PsyMem (Role-based PD data) 81.13 85.20 89.93 95.60
Qwen2.5-7B-Instruct 79.14 64.40 89.58 97.40
     + PsyMem (Original PD data) 81.63 86.60 89.45 93.40
     + PsyMem (Role-based PD data) 82.64 87.60 89.64 96.20
Table 5: Ablation study on Role-based (RB) Pure Dove (PD) dataset in the PsyMem framework.
Figure 4: The impact of data size on model performance using the Pure Dove (PD) dataset.
We further analyzed the effectiveness of memory injection and the incorporation of role-based general SFT data in training.
Memory alignment training. To evaluate the impact of memory alignment training on character fidelity, we constructed a memory testing dataset using GPT-4o, designing 1,500 questions across 50 roles in 500 scenarios based on detailed character profiles. We present our results in Table 3, comparing the performance of PsyMem-LLama and PsyMem-Qwen, with and without memory alignment training. Our findings indicate that incorporating memory data in the second training stage significantly enhances the model’s ability to responding based on retrieved memory, rather than merely summarizing relevant content. This improves both character attribute consistency and long-term memory retention. However, we also observe that memory alignment training has a slight negative impact on character attribute alignment. We hypothesize that this may be due to the presence of a small number of role-specific dialogue examples in the memory data, which could slightly weaken the model’s ability to strictly align with predefined character attributes.
Role-Based General SFT Data We investigate the impact of Role-based General SFT Data on Character-Independent Capabilities and Character Fidelity in Table 5 and Figure 4. Table 5 presents the ablation experiment on the use of Role-based General SFT Data, showing that PsyMem-LLama and PsyMem-Qwen helps mitigate the decline in coherence capabilities caused by catastrophic forgetting of general capabilities. Additionally, it leads to improvements in Character Fidelity, Human-likeness, and Consistency.
In Figure 4, we examine the impact of General SFT Data and Role-based General SFT Data on Character Fidelity for PsyMem-Qwen. The results validate the effectiveness of Role-based General SFT Data. Specifically, we observe that the use of General SFT Data negatively impacts the model’s character-specific abilities, whereas training with Role-based General SFT Data achieves optimal performance with 1,500 samples, after which performance begins to decline. Despite this, it consistently outperforms General SFT Data. We hypothesize that the decline in performance with 3,000 samples of Role-based General SFT Data occurs because the dataset’s distribution becomes overly dominated by GPT-generated content, causing an imbalance in role-specific attributes. A balanced amount of Role-based General SFT Data thus helps improve character-playing abilities without leading to such issues.
6 Conclusion
In this work, we propose a novel framework PsyMem for controlling character role-playing in large language models (LLMs) by integrating two key aspects: character memory and character attributes, both of which align more closely with modern psychological theories. We validate the proposed PsyMem using a large scale specially designed role-playing dataset based on real novels, addressing the gap between generic and well-known characters in role-playing tasks. Our experiments show that PsyMem-Qwen, trained with our alignment across five dimensions, outperforms all baselines mentioned in this paper, demonstrating strong performance in Character-Independent Capabilities as well.
Furthermore, we explore the benefits of incorporating Role-based General SFT Data for role-playing models. Through this work, we aim to encourage future research on role-playing models that consider psychological frameworks for a more nuanced understanding of character behavior, ultimately advancing the development of more realistic and consistent role-playing LLMs.

### v2

@@
 Recent advancements in large language models (LLMs) (Achiam et al., 2023; Anthropic, 2024a; Anil et al., 2023) have unlocked new possibilities for implementing role-playing systems through language-based behavior simulation (Shanahan et al., 2023).
-Role-playing systems demonstrate societal value across applications like interactive gaming (Wang et al., 2023a) and AI counseling (Zhang et al., 2024), particularly in enabling social simulation.
-Such simulation allows modeling of human interactions and societal dynamics (Park et al., 2023), offering insights into policy design and conflict mediation (Zhou* et al., 2024).
+Role-playing systems demonstrate societal value across applications like interactive gaming (Wang et al., 2023a) and AI counseling (Zhang et al., 2024), with particularly potential in enabling high-reliable social simulation.
+Such simulation allows modeling of human interactions and societal dynamics (Park et al., 2023), offering insights into evidence-based policy design and conflict mediation (Zhou* et al., 2024).
 To achieve trustworthy social simulation, LLM-based role-playing must preserve persistent traits, memories, and behavioral logic to ensure agents reliably mirror specified attributes in evolving social contexts.
@@
 In this work, we propose a widely recognized modern psychological evaluation framework (Costa and McCrae, 2008; Schwartz, 1992; Thomas, 2008; Zimbardo and Ruch, 1975) and a graph-structured memory mechanism to construct the PsyMem framework, enabling LLMs to achieve exceptional role-playing abilities.
-Drawing from modern psychology and neuroscience, we enhance the description of character attributes by leveraging both latent psychological attributes and explicit behavioral patterns (Snyder, 1983; Kahneman, 2011), with personality and values corresponding to latent psychological attributes, and behavioral decision-making, social interaction, and leadership representing explicit behavioral patterns.
-Overall, we describe characters using 26 quantitative indicators and a small amount of textual descriptions.
-We extract real data from novels and construct a publicly available role-playing dataset.
+Drawing from modern psychology and neuroscience, we enhance the description of character attributes by leveraging both latent psychological attributes and explicit behavioral patterns (Snyder, 1983; Kahneman, 2011), with personality and values corresponding to latent psychological attributes, and behavioral decision, social interaction, and leadership representing explicit behavioral patterns.
+Overall, we extract real data from novels and construct a large scale role-playing dataset, using 26 quantitative indicators and a small amount of textual descriptions to describe characters.
 As shown in Table 1, this dataset contains 5,414 established characters, 38,962 conversations, and 536,636 utterances, much larger than existing role-playing datasets.
@@
   • We propose PsyMem, a novel LLM-based role-playing framework that establishes precise control over character behaviors through the synergistic integration of fine-grained psychological profiling (e.g., personality, value) and dynamic memory schemata.
-  • Beyond textual descriptions, we present a comprehensive set of attributes that bridges intrinsic (e.g., values, emotions) and extrinsic (e.g., behaviors, social interactions) character dimensions, operationalized via 26 measurable indicators.
+  • Beyond textual descriptions, we present a comprehensive attribute set that bridges intrinsic (e.g., values, emotions) and extrinsic (e.g., behaviors, social interactions) character dimensions, operationalized via 26 measurable indicators.
 This attribute set captures both the internal motivations and external behavioral manifestations of characters.
   • We introduce a memory alignment training paradigm that establishes explicit cognitive anchoring between role-playing responses and character memory.
-This methodology enhances role-playing through a dual-control mechanism, enabling precise regulation of generative outputs via both trait-based attributes (e.g., personality dimensions) and episodic memory constraints.
-  • We introduce the use of role-based general supervised fine-tuning data (transforming general SFT data into role-playing style SFT data), demonstrating that incorporating such data effectively enhances the performance of Role-Playing LLMs.
+This methodology enhances role-playing through a dual-control mechanism, enabling precise regulation of generative outputs via both attributes (e.g., personality dimensions) and episodic memory constraints.
+  • We introduce the use of role-playing style general supervised fine-tuning data (transforming general SFT data into role-playing style), demonstrating that incorporating such data effectively enhances the performance of Role-Playing LLMs.
   • Extensive experiments demonstrate that the proposed PsyMem framework significantly improves the role-playing ability of LLMs.
@@
 We argue that a character’s behavior is jointly influenced by their attributes, memory, and the scenario.
-Drawing from modern psychology, we segment character attributes using contemporary psychological frameworks and construct a dataset containing 38,962 conversations (536,636 utterances) from real novels, leveraging both character attributes and memory to enhance the model’s role-playing ability.
+Drawing from modern psychology, we segment character attributes using contemporary psychological frameworks and construct a dataset containing 38,962 conversations (536,636 utterances) from real novels, leveraging both character attributes and memory to enhance role-playing ability.
 2.2 Memory Mechanism in Role-Playing.
@@
 For example, Shao et al.
-(2023) and others push the boundaries of knowledge to protect character memory, while Lu et al.
+(2023) and others Lu et al.
+(2024); Yu et al.
+(2024) push the boundaries of knowledge to protect character memory, while Lu et al.
 (2024) and others use a self-alignment approach to distinguish the different memory backgrounds of various characters.
-Relying on the internal knowledge within the model is effective for famous characters but fails for infamous and unseen characters.
-Some works apply Retrieval-Augmented Generation (RAG) (Ram et al., 2023) which dynamically retrieving external contexts to explicitly mimic character’s memory.
-For example, Zhong et al.
-(2024) adopts a memory update mechanism inspired by Ebbinghaus’s forgetting curve theory, emphasizing the recall of relevant memories to understand the user’s personality.
+Relying on the internal knowledge within the model is effective for famous characters but fails for unfamous and unseen characters.
+Some works apply Retrieval-Augmented Generation (RAG) (Ram et al., 2023) to dynamically retrieve external contexts for explicitly mimicking character memory, such as Zhong et al.
+(2024) who adopts memory update mechanisms based on Ebbinghaus’s forgetting curve theory.
 However, as shown in Table 2, existing role-playing models tend to perform better when simulating famous characters, while their performance significantly drops for ordinary characters that rely on external memory.
@@
 In contrast, “LLMs as Judges”, with its efficiency, low cost, and strong scalability, is increasingly becoming the preferred choice.
+The experiments in Appendix (see Table 7) demonstrate the effectiveness of the “LLMs as Judges” paradigm for role-playing evaluation.
 3 Dataset Architecture
@@
 Moreover, existing datasets predominantly focus on character attributes, overlooking the significant influence memory has on characters’ behaviors, thought processes, and speech styles(Stumpf and Dunbar, 1991; Ravlin and Meglino, 1987; Pickering and Garrod, 2004).
-This limitation becomes particularly evident when role-playing characters possess limited background knowledge during the pretraining stage (see Figure 1).
+This limitation becomes particularly evident when LLMs can hardly access the background knowledge of the role-playing characters during the pretraining stage (see Figure 1).
 To address these shortcomings, we design a novel dataset with the architecture formulated as
@@
 where $N$ is the number of data sample.
-Each sample is a dialogue turn where $R_{i}$ is the role for response; $P_{i}$ is the profile of $R_{i}$, featuring 26 quantitative psychological dimensions; $C_{i}$ is the previous dialogue turns before the current dialogue.
+Each sample is a real dialogue turn collected from novels.
+$R_{i}$ is the role for response; $P_{i}$ is the profile of $R_{i}$, including basic information and 26 quantitative psychological dimensions; $C_{i}$ consists of the dialogue turns preceding the current turn, as well as the scene information.
 $x_{i}$ is the query to $R_{i}$ and $y_{i}$ is the corresponding response.
-$M_{i}$ is the retrieved memory response of role $R_{i}$ with respect to query $x_{i}$ and the previous dialogue $C_{i}$.
-In this work, the dataset $D_{\mathrm{RP}}$ contains 38,962 samples, extracted from 539 novels characterized by a wide range of genres, rich content, and narrative perspectives of third person.(The genre distribution is presented in Figure 2.)
+$M_{i}$ denotes the memory retrieved for role $R_{i}$, conditioned on the query $x_{i}$ and the context $C_{i}$, which comprises the preceding dialogue turns and the scene.
+In this work, the dataset $\text{D}_{\text{RP}}$ contains 38,962 samples, extracted from 539 novels characterized by a wide range of genres, rich content, and narrative perspectives of third person.(The genre distribution is presented in Figure 2.)
 Dataset Character Conversation Profile Style
@@
 These values deeply affect decision-making, priorities, ethical judgments, and interpersonal alignment, enriching the psychological depth and motivational dimensions of individuals.
-Explicit behavioral patterns manifest externally in observable actions, interactions, and communication styles. We capture these explicit patterns using:
+Explicit behavioral patterns manifest in observable actions, interactions, and communication styles. We capture these explicit patterns using:
 1) Social & Leadership Styles: Measured through six dimensions based on Zimbardo’s psychological principles (Zimbardo and Ruch, 1975), significantly influencing observable interpersonal behaviors, communication effectiveness, leadership approaches, and social dynamics in groups.
@@
 Therefore, we introduce character memory data $M$ to the role playing dataset to enable the memory simulation capability of role playing model.
-Specifically, we first build a knowledge graph based on the novel via GraphRAG (Edge et al., 2024), and then retrieve relevant memory for the current dialogue via entity mentions, relational triples, and raw dialogue chunks associated with specific scenes.
+Specifically, we first build a knowledge graph for each novel via GraphRAG (Edge et al., 2024), and then retrieve relevant memory context for the current dialogue in local mode (see Appendix B.1 for details).
 Memory alignment is evaluated on two key dimensions: correctness and rationality.
@@
 These were integrated into novel-based dialogues to maintain coherence and character fidelity.
-3.3 Role-based SFT Data
-We propose a Role-based general supervised fine-tuning (SFT) data construction approach.
-Dialogue samples from the Pure-Dove (PD) dataset (Daniele and Suphavadeeprasit, 2023) are grouped randomly, and LLMs infer each speaker’s latent psychological attributes and explicit behavioral patterns.
-Scene-specific information is dynamically generated using fine-tuned LLMs based on dialogue contexts, significantly improving character consistency and alignment during interactions, as demonstrated by our experimental evaluations.
-Similar to Equation 1, we construct a dataset defined as:
+3.3 Synthetic Role-playing Data
+In addition to $\text{D}_{\text{RP}}$, we propose a synthetic role-playing data construction approach that transforms general supervised fine-tuning (SFT) data into role-playing style dialogues, resulting in the dataset $\text{D}_{\text{RP}}^{\text{synth}}$.
+The construction pipeline consists of three steps: 1) We randomly sample multi-turn dialogues from the Pure-Dove (PD) dataset (Daniele and Suphavadeeprasit, 2023) and sequentially group every 100 consecutive dialogue turns to form speaker-specific clusters.
+Each cluster containing approximately 100 dialogue turns and is then assigned to represent a distinct synthetic speaker.
+2) For each synthetic speaker, we use GPT-4o to analyze the 100 dialogue turns and annotate the speaker’s psychological profile according to our 26 quantitative dimensions.
+3) Based on the dialogue context, we employ GPT-4o to generate appropriate scene-specific information to enhance contextual coherence.
+The resulting dataset $\text{D}_{\text{RP}}^{\text{synth}}$ is formulated as:
 […]
-Where $R_{i}$ denotes the role responsible for the response, and $P_{i}$ is the profile of $R_{i}$ defined by 26 quantitative psychological dimensions.
-$C_{i}$ represents the preceding dialogue context, $x_{i}$ is the current query directed to $R_{i}$, and $y_{i}$ is the corresponding response.
-These $(x_{i},C_{i},y_{i})$ triples are drawn from general supervised fine-tuning datasets.
-In addition, $M_{i}$ refers to the retrieved memory response relevant to role $R_{i}$ with respect to $x_{i}$ and $C_{i}$.
+where $R_{i}$ denotes the role responsible for the response, and $P_{i}$ is the profile of $R_{i}$ defined by basic information and 26 quantitative psychological dimensions.
+$C_{i}$ represents the preceding turns and scene information.
+$x_{i}$ and $y_{i}$ are the current query, and the corresponding response drawn from the Pure-Dove (PD) datasets, respectively.
+By incorporating role-playing style general SFT data $\text{D}_{\text{RP}}^{\text{synth}}$, the proposed PsyMem framework enables the model to develop stronger role-playing capabilities, as Figure 4 shows.
 4 Role Playing Training and Evaluation
@@
 This challenge exemplifies the fundamental tension in transfer learning between preserving general knowledge and specializing for domain-specific tasks.
-Traditional fine-tuning approaches (Kotha et al., 2024) often suffer from catastrophic forgetting, where optimization for character fidelity compromises the model’s general capabilities.
-Inspired by the Dual-stage Mixed Fine-tuning (DMT) (Dong et al., 2023) method, we introduce a two-stage training strategy for role-playing, effectively balance character fidelity and the preservation of general language understanding capabilities, as illustrated in Figure 3b.
-Specifically, we introduce a two-stage progressive training framework to systematically develop LLMs into specialized Role-Playing Models: Stage 1 establishes core role-playing capabilities through fine-tuning on data emphasizing fundamental role characteristics.
-Stage 2 enhances specialization via memory-augmented fine-tuning, incorporating long-term memory contexts and general role-based information.
-This approach ensures a balanced development of character-specific attributes while simultaneously preserving the model’s general capabilities.
-Stage 1: Foundational Role-Playing Capacity Development.
-In this stage, we train LLMs to learn foundational role-playing ability based only on quantifiable attributes and basic information.
-We fine-tune the model using a subset $D_{\mathrm{RP1}}$ sampled from the overall role-playing dataset to establish its fundamental role-playing abilities.
+Traditional fine-tuning approaches (Kotha et al., 2024) often encounter a core multi-task conflict during optimization.
+In the context of role-playing tasks, this conflict arises from the attempt to optimize two fundamentally distinct objectives—namely, character profile alignment and memory adherence—within a unified framework.
+Inspired by the Dual-stage Mixed Fine-tuning (DMT) (Dong et al., 2023) method, we introduce a two-stage training strategy for role-playing, effectively balancing the capabilities of character profile alignment and memory adherence.
+Specifically, stage 1 establishes core role-playing capabilities by training the model on character profile-based dialogues, enabling the model to learn fundamental character attribute alignment.
+Stage 2 further trains the model on memory-augmented role-playing dataset $\text{D}_{\text{RP}}$ and role-playing style general fine-tuning dataset $\text{D}_{\text{RP}}^{\text{synth}}$ to improve role-playing precision, as illustrated in Figure 3(b).
+As demonstrated in Section 5.3, our two-stage training strategy significantly outperforms vanilla single-stage training for role-playing performance, validating the necessity of this progressive specialization strategy.
+Stage 1: Foundational Role-Playing Capacity Development
+In this stage,we train LLMs to learn foundational role-playing abilities based on quantifiable character attributes and dialogue contexts without memory augmentation.
+We fine-tune the model using a subset $\text{D}_{\text{RP1}}$ sampled from the overall role-playing dataset $\text{D}_{\text{RP}}$ to establish its fundamental role-playing abilities.
 The objective function is formally defined as:
 […]
-where ()($\cdot$)( ⋅ ) in $D_{\mathrm{RP}}$ represents the tuple $(R,P,C,x,y)$, $D_{\mathrm{RP1}}$ is a subset sampled from the overall role-playing dataset $D_{\mathrm{RP}}$, excluding the retrieved memory $M$.
-Stage 2: Integrated Memory-Augmented Specialization.
-To achieve integrated memory-augmented specialization, we further specialize the model by fine-tuning it on a distinct subset from the role-playing dataset $D_{\mathrm{RP}}$, augmented with long-term memory contexts retrieved via graph-based methods, and combined with general role-based SFT data $D_{\mathrm{RB}}$.
-This stage employs a weighted combination of two loss terms, explicitly designed to mitigate forgetting of general capabilities while simultaneously enhancing the model’s ability to effectively leverage long-term memory contexts:
-[…]
+where $(\cdot)$ in $\text{D}_{\text{RP1}}$ denotes the tuple $(R,P,C,x,y)$ and $\text{D}_{\text{RP1}}$ excludes the retrieved memory component $M$ from the original dataset $\text{D}_{\text{RP}}$.
+Stage 2: Memory-Augmented Role-Playing Specialization.
+In this stage, we enhance the model’s role-playing capabilities through two training components: 1) memory-augmented role-playing data $\text{D}_{\text{RP}}$, which incorporates long-term memory contexts retrieved via graph-based methods to enable explicit memory alignment, and (2) synthetic role-playing style data $\text{D}_{\text{RP}}^{\text{synth}}$, which improves general role-playing abilities.
+This stage employs a weighted combination of two corresponding loss terms to simultaneously enhance memory-controlled response generation while preserving the model’s foundational capabilities:
+$\theta^{*}_{2}=\arg\min_{\theta}\left\{\alpha\mathcal{L}_{\mathrm{RP}}(\theta)+\mathcal{L}_{\mathrm{RP}}^{\mathrm{synth}}(\theta)\right\}$
 where $\alpha=20$.
@@
 […]
-[…]
-where ()($\cdot$)( ⋅ ) in $D_{\mathrm{RP2}}$ represents the tuple $(R,P,M,C,x,y)$ and in $D_{\mathrm{RB}}$ represents $(R,P,C,x,y)$.
-$M$ denotes the retrieved memory response for role $R$ regarding the query $x$ and prior dialogue context $C$, obtained via graph-based retrieval methods and subsequently filtered by LLMs.
+$\mathcal{L}_{\mathrm{RP}}^{\mathrm{synth}}(\theta)=\mathbb{E}_{(\cdot)\sim\text{D}_{\text{RP}}^{\text{synth}}}\left[\mathcal{L}\left(f_{\theta}(R,P,C,x),y\right)\right]$
+where $(\cdot)$ in $\text{D}_{\text{RP}}^{\text{synth}}$ represents $(R,P,C,x,y)$ and $(\cdot)$ in $\text{D}_{\text{RP}}$ represents the tuple $(R,P,M,C,x,y)$.
+$M$ denotes the retrieved memory response for role $R$ regarding the query $x$ and the context $C$ (which includes the preceding dialogue turns and scene information), obtained via graph-based retrieval methods and subsequently filtered by LLMs.
 Model Character Fidelity Character-independent
@@
 4.2 Evaluation
-We evaluate role-playing capabilities of LLMs by adopting the widely used “LLMs as Judges” approach (Shao et al., 2023; Lu et al., 2024), as illustrated in Figure 3c.
+We evaluate role-playing capabilities of LLMs by adopting the widely used “LLMs as Judges” approach***We validate the reliability of our LLM-based evaluation framework by comparing human-evaluation and GPT-4o judgments in Appendix (see Table 7) (Shao et al., 2023; Lu et al., 2024), as illustrated in Figure 3(c).
 The evaluation process is structured into two main categories: Character-independent Capabilities and Character Fidelity.
-Additionally, in the Tabel 4, we further discuss the discrepancies between LLM-based judgments and human evaluations across 26 quantitative indicators to validate the effectiveness of the “LLMs as Judges” approach.
-4.2.1 Character-independent Capabilities
+Character-independent Capabilities.
 Following previous research (Chen et al., 2024; Yu et al., 2024) , we measure Character-independent Capabilities using three metrics: Human-likeness, Consistency and Coherence.
@@
 For consistency, we evaluate the coherence between different parts of the conversation to assess whether the role-playing remains consistent throughout long dialogues.
-Additionally, for coherence, we assess how logically connected and contextually appropriate the responses are across the entire dialogue, ensuring the conversation flows naturally.
-4.2.2 Character Fidelity
+For coherence, we assess how logically connected and contextually appropriate the responses are across the entire dialogue, ensuring the conversation flows naturally.
+Character Fidelity.
 Character Fidelity assesses how accurately an LLM portrays specific role-playing characters.
-As described in Section 3.1, we employ modern psychological quantification frameworks defined in our dataset design, covering four primary dimensions: Personality (Big Five Personality Model), Values (Schwartz’s Theory of Basic Values), Social and Leadership Styles (Zimbardo’s psychological principles), and Behavioral Decision-making (Thomas-Kilmann Conflict Mode Instrument, TKI).
-These dimensions are quantified into numeric scales using standardized psychological evaluation methods.
-Additionally, we separately evaluate the model’s memory, focusing on the correctness and rationality of memory alignment.
-The Memory dimension is measured using accuracy scores derived from a true/false evaluation format.
-To calculate similarity between role-playing characters, we employ Normalized Euclidean Distance (NED), computed as follows:
+As described in Section 3.1, we employ modern psychological quantification frameworks defined in our dataset design, covering five primary dimensions: Personality (Big Five Personality Model), Values (Schwartz’s Theory of Basic Values), Social and Leadership Styles (Zimbardo’s psychological principles), Behavioral Decision (Thomas-Kilmann Conflict Mode Instrument, TKI), and Memory.
+The former four dimensions are quantified into numeric scales ranging from 1 to 5 using standardized psychological evaluation methods.
+The Memory dimension, distinct from the psychological attributes, is evaluated using accuracy scores from true/false assessments.
+Evaluation Process.
+The evaluation process commences with models generating 15-turn dialogues per scenario (see Section 5.1 for scenario detail), embodying the specified character profile.
+For character-independent capabilities evaluation, we utilize these generated dialogues: 1) Human-likeness and Coherence are assessed by GPT-4o (true/false based on rules), yielding a positive assessment rate (e.g., the percentage deemed human-like or coherent).
+2) Consistency is measured by the score difference in the four psychological dimensions between dialogue halves, yielding a stability score (e.g., where a smaller difference yields a higher consistency score, scaled from 0 to 1).
+For character fidelity evaluation, GPT-4o serves as the judge, assessing each 15-turn dialogue.
+Performance is scored across five dimensions: 1) The first four dimensions (Personality, Values, etc.,) are evaluated by GPT-4o on dialogues, using 1-NED (introduced in the following paragraph) for performance; 2) For the fifth dimension, memory alignment, we use GPT-4o to judge response correctness (true/false) against memory context and reference answer, yielding an accuracy score (i.e., the proportion of correctly aligned responses).
+Metrics.
+We quantify role-playing fidelity for Personality, Values, Social & Leadership, and Behavioral Decision using the 1-NED metric.
+Derived from Normalized Euclidean Distance (NED), a higher 1-NED score indicates greater similarity to the target profile.
+The 1-NED score is computed as:
 […]
-where $\Delta X=X_{\max}-X_{\min}$, $n$ is the number of subdimensions, $X_{\mathrm{ref}}$ is the standardized reference value, and $X_{i}$ represents the measurement for the $i$-th subdimension, with possible values from $Z=\{1,2,3,4,5\}$.
+where $\Delta\!X=\!\!X_{\max}\!\!-\!X_{\min}$, with $X_{\max}$ and $X_{\min}$ being the maximum (5) and minimum (1) values of the evaluation scale, respectively.
+A score of 0, used to denote uncertainty, is excluded from this range.
+$n$ is the number of subdimensions, $X_{i}$ represents the measurement for the $i$-th subdimension, and $X_{\mathrm{ref}}$ is the corresponding value contained in role profile $P_{i}$ (see Eq.(1) and Eq.(2)), with possible values from $\{1,2,3,4,5\}$.
 This formulation is robust, as it normalizes the mean squared deviation across different scales, yielding a more interpretable similarity metric.
 5 Experiments
-In this section, we explain the dataset, baselines, and metrics design in the experimental setup, and discuss the main results. Additional ablation experiments are discussed in Section 5.3.
+In this section, we explain the dataset, baselines, and implementation, and discuss the main results and ablation experiments.
 5.1 Experimental Setup
 Evaluation Dataset.
-To mitigate potential knowledge bias introduced during the pretraining phase, we construct an evaluation dataset for our main experiments by selecting characters and scenes from novels published after June 2024.
+To mitigate potential knowledge bias introduced during the pretraining phase, we construct an evaluation dataset $\text{D}_{\text{RP}}^{\text{eval}}$ for our main experiments by selecting characters and scenes from novels published after June 2024.
 This design choice aims to ensure that models have had minimal or no exposure to the selected texts during pretraining, thereby reducing reliance on memorized character-specific knowledge.
-We first score 100 contemporary novels using large language models (LLMs) and then select 50 representative characters spanning diverse genres.
-For each character, we collect ten scenes directly sourced from the original text.
+We first assess 100 contemporary novels across diverse genres, then select 25 high-quality novels based on character complexity, dialogue richness, expression quality, and narrative perspective.
+From each selected novel, we identify the two most psychologically complex and representative characters, yielding 50 characters with varied psychological profiles.
+For each character, we extract ten dialogue-rich scenes directly from the original novels, resulting in 500 evaluation scenarios (50 characters × 10 scenes), which constitutes our evaluation set $\text{D}_{\text{RP}}^{\text{eval}}$.
+The character attributes are derived from a detailed analysis of the novel using GPT-4o.
+This comprehensive evaluation framework enables robust assessment of role-playing capabilities across diverse character types and narrative contexts.
+For the memory evaluation, each scenario includes not only the character’s profile and scene information, but also three key components: a meticulously selected character memory excerpt from the memory text, a memory-dependent question, and its authentic answer serving as the reference answer, allowing us to assess all character fidelity dimensions.
 Baselines.
-In our benchmark evaluation, we tested both open-source and proprietary advanced chatbots, each representing the state-of-the-art within their respective frameworks.
+Baselines in our evaluation include both open-source and proprietary advanced chatbots, each representing the state-of-the-art within their respective frameworks.
 For general-purpose models, we compared the proposed PsyMem with several prominent baselines, including GPT-4o (Hurst et al., 2024), GPT-3.5-turbo (OpenAI, 2023),Claude-3.5-sonnet (Anthropic, 2024b), Qwen-Max (Team, 2024), Yi-Large (Young et al., 2024), Deepseek-R1 (Guo et al., 2025), DeepSeek-V3 (Liu et al., 2024), LLaMA3.1-8B-instruct (Grattafiori et al., 2024) and Qwen2.5-7B-Instruct (Yang et al., 2024).
@@
 Implementations.
-With the proposed PsyMem framework, we train our models based on Llama3.1-8B-Instruct and Qwen2.5-7B-Instruct, and denote the trained models as PsyMem-LLama and PsyMem-Qwen, respectively.
-All models were trained on four A800 80G GPUs for 3 epochs with a global batch size of 128 and a sequence length of 8,192.
-LoRA was used with a per-device batch size of 2 and 16 gradient accumulation steps.
-We adopted a cosine learning rate schedule with a warmup ratio of 0.1, and set the learning rate to 1e-4.
-Metrics
-We utilize GPT-4o to evaluate 50 established characters across 500 scenarios, each scenario comprising 15 dialogue turns.
-Evaluations are conducted based on the dimensions described in Section 3, with mean scores calculated accordingly.
+With the proposed PsyMem framework, we fine-tune LLaMA3.1-8B-Instruct and Qwen2.5-7B-Instruct for three epochs, and denote the resulting models as PsyMem-LLaMA and PsyMem-Qwen, respectively.
+Additional training settings and implementation details are provided in the Appendix B.3.
+Evaluations are conducted according to Section 4.2.
 5.2 Main Results
 We present the main results in Table 2.
-We extracted 500 real novel scenes to evaluate all models, assessing them from two perspectives: Character Fidelity and Character-independent Capabilities.
-Notably, in terms of Memory alignment, Proprietary models demonstrate a more reasonable approach—not merely summarizing but leveraging memory to portray characters more authentically.
-Although Higgs-LLaMA3-70B and CoSER-70B exhibit relatively stronger performance in memory handling compared to other role-play-specialized models, there remains a noticeable gap between them and top-performing general-purpose models.
-Within the General Baselines, we observe that Proprietary models still significantly outperform open-source models.
-GPT-4o and DeepSeek-V3, as outstanding proprietary models, exhibit impressive performance in Character-independent Capabilities.
-Additionally, Baichuan-NPC-Turbo, as a role-play expertise baseline model, surpasses general baselines in self-consistency, indicating that this metric is particularly important for role-playing agents.
+Within the General Baselines, we observe that proprietary models commonly outperform open-source models.
+Claude-3.5-sonnet and Yi-Large, as outstanding proprietary models, exhibit impressive performance in Character-independent Capabilities.
+Additionally, Baichuan-NPC-Turbo, as a role-play expertise baseline model, surpasses general baselines in self-consistency, which is particularly important for role-playing agents.
 Furthermore, it is worth mentioning that DeepSeek-R1 and CoSER-70B excel in Character Fidelity.
-Empirical results under the PsyMem framework reveal that, relative to their original counterparts, PsyMem-Qwen and PsyMem-LLaMA achieve character fidelity improvements of 2.07% and 3.50%, respectively.
+Empirical results under the PsyMem framework reveal that, relative to their original counterparts, PsyMem-LLaMA and PsyMem-Qwen achieve character fidelity improvements of 2.07% and 3.50%, respectively.
 In addition, PsyMem-Qwen demonstrates strong performance in the Social & Leadership and Behavioral Decision dimensions, achieving scores of 80.80% and 78.48%, respectively.
@@
 Model Character Fidelity
-Per. Val. SL* BD* Mem. Avg.
+Per. Val. SL BD Mem. Avg.
 LLaMA3.1-8B-Instruct
 + PsyMem (w/o Mem.) 80.11 81.90 79.23 76.36 44.20 72.36
++ PsyMem (w/o DMT.) 79.97 81.34 77.96 74.55 89.40 80.64
 + PsyMem 80.47 81.93 78.57 74.47 90.20 81.13
@@
 + PsyMem (w/o Mem.) 81.02 81.33 82.10 79.02 53.00 75.29
++ PsyMem (w/o DMT.) 79.59 81.55 80.37 77.96 90.40 81.97
 + PsyMem 80.40 81.74 80.80 78.48 91.80 82.64
-Table 3: Ablation study on memory integration in the PsyMem framework. Per.: Personality, Val.: Values, SL: Social & Leadership, BD: Behavioral Decision, Mem.: Memory.
-Metrics Per. Val. SL* BD*
-Cosine 0.98 0.96 0.97 0.97
-Table 4: Cosine similarity between human and GPT-4o scores. Per.: Personality, Val.: Values, SL: Social & Leadership, BD: Behavioral Decision.
+Table 3: Ablation studies on DMT. and memory integration within the PsyMem framework.
 5.3 Ablation Study
@@
 LLaMA3.1-8B-Instruct 79.06 64.20 88.74 99.00
-     + PsyMem (Original PD data) 80.51 70.80 89.49 82.20
-     + PsyMem (Role-based PD data) 81.13 85.20 89.93 95.60
+     + PsyMem (PD data) 80.51 70.80 89.49 82.20
+     + PsyMem ($\text{D}_{\text{RP}}^{\text{synth}}$) 81.13 85.20 89.93 95.60
 Qwen2.5-7B-Instruct 79.14 64.40 89.58 97.40
-     + PsyMem (Original PD data) 81.63 86.60 89.45 93.40
-     + PsyMem (Role-based PD data) 82.64 87.60 89.64 96.20
-Table 5: Ablation study on Role-based (RB) Pure Dove (PD) dataset in the PsyMem framework.
-Figure 4: The impact of data size on model performance using the Pure Dove (PD) dataset.
-We further analyzed the effectiveness of memory injection and the incorporation of role-based general SFT data in training.
+     + PsyMem (PD data) 81.63 86.60 89.45 93.40
+     + PsyMem ($\text{D}_{\text{RP}}^{\text{synth}}$) 82.64 87.60 89.64 96.20
+Table 4: Ablation study on Synthetic Role-playing style ($\text{D}_{\text{RP}}^{\text{synth}}$) Pure Dove (PD) dataset in PsyMem.
+Figure 4: Comparing Original and Synthetic Role-Playing data ($\text{D}_{\text{RP}}^{\text{synth}}$). PD: Performance by Dataset Size.
 Memory alignment training.
-To evaluate the impact of memory alignment training on character fidelity, we constructed a memory testing dataset using GPT-4o, designing 1,500 questions across 50 roles in 500 scenarios based on detailed character profiles.
-We present our results in Table 3, comparing the performance of PsyMem-LLama and PsyMem-Qwen, with and without memory alignment training.
+To evaluate the effectiveness of memory alignment training, we compare the performance of PsyMem-LLaMA and PsyMem-Qwen with and without memory integration during the second training stage.
+The results are presented in Table 3.
 Our findings indicate that incorporating memory data in the second training stage significantly enhances the model’s ability to responding based on retrieved memory, rather than merely summarizing relevant content.
@@
 However, we also observe that memory alignment training has a slight negative impact on character attribute alignment.
-We hypothesize that this may be due to the presence of a small number of role-specific dialogue examples in the memory data, which could slightly weaken the model’s ability to strictly align with predefined character attributes.
-Role-Based General SFT Data We investigate the impact of Role-based General SFT Data on Character-Independent Capabilities and Character Fidelity in Table 5 and Figure 4.
-Table 5 presents the ablation experiment on the use of Role-based General SFT Data, showing that PsyMem-LLama and PsyMem-Qwen helps mitigate the decline in coherence capabilities caused by catastrophic forgetting of general capabilities.
-Additionally, it leads to improvements in Character Fidelity, Human-likeness, and Consistency.
-In Figure 4, we examine the impact of General SFT Data and Role-based General SFT Data on Character Fidelity for PsyMem-Qwen.
-The results validate the effectiveness of Role-based General SFT Data.
-Specifically, we observe that the use of General SFT Data negatively impacts the model’s character-specific abilities, whereas training with Role-based General SFT Data achieves optimal performance with 1,500 samples, after which performance begins to decline.
-Despite this, it consistently outperforms General SFT Data.
-We hypothesize that the decline in performance with 3,000 samples of Role-based General SFT Data occurs because the dataset’s distribution becomes overly dominated by GPT-generated content, causing an imbalance in role-specific attributes.
-A balanced amount of Role-based General SFT Data thus helps improve character-playing abilities without leading to such issues.
+We hypothesize that this may be due to the inclusion of a limited number of role-specific dialogue examples in the memory data, which could slightly diminish the model’s ability to adhere strictly to predefined character attributes.
+Two-stage Training Strategy.
+To evaluate the effectiveness of the DMT strategy, which serves as the core of our two-stage training strategy approach, we conducted ablation studies comparing the full PsyMem framework (with DMT) against a single-stage fine-tuning baseline, in which all datasets and loss functions are jointly optimized.
+As shown in Table 3, DMT consistently improves performance across both backbone models (Qwen and LLaMA).
+For PsyMem-Qwen, DMT leads to an overall improvement of 0.67% (82.64% vs.
+81.97%), while for PsyMem-LLaMA, the improvement is 0.49% points (81.13% vs.
+80.64%), demonstrating the effectiveness of the two-stage approach in enhancing character fidelity.
+Synthetic Role-playing Data.
+We investigate the impact of Synthetic Role-playing style SFT Data $\text{D}_{\text{RP}}^{\text{synth}}$ on model performance through ablation studies presented in Table 4 and Figure 4.
+Table 4 compares model performance using $\text{D}_{\text{RP}}^{\text{synth}}$ versus original general SFT data.
+The results demonstrate that incorporating $\text{D}_{\text{RP}}^{\text{synth}}$ significantly improves Character Fidelity, Human-likeness, and Consistency for both PsyMem-LLaMA and PsyMem-Qwen, while effectively maintaining coherence capabilities.
+Figure 4 further examines the effect of data volume, comparing General SFT Data with $\text{D}_{\text{RP}}^{\text{synth}}$ on role-playing performance.
+The results validate the superiority of our synthetic approach: while General SFT Data consistently degrades character-specific abilities, Synthetic Role-playing style data achieves optimal performance at 1,500 samples before declining at larger volumes.
+We attribute this performance drop at 3,000 samples to dataset distribution imbalance, where excessive synthetic content may dilute role-specific attribute diversity.
+Figure 5: Ablation study of each dimensions. Per.: Personality, Val.: Values, SL: Social & Leadership, BD: Behavioral Decision, Mem.: Memory.
+Character Attribute Dimension.
+We assess the contribution of each character attribute by conducting an ablation study on PsyMem-Qwen, where each of the four supervision signals—Personality, Values, Social & Leadership, and Behavioral Decision—is individually removed (see Figure 5).
+The results reveal that removing any single dimension markedly drops performance on its corresponding evaluation metric.
+This indicates each dimension is critical to the model’s fidelity for that trait, underscoring the necessity of the proposed comprehensive attribute set.
+We have placed more ablation experiments on dimensions in Appendix D.
 6 Conclusion
-In this work, we propose a novel framework PsyMem for controlling character role-playing in large language models (LLMs) by integrating two key aspects: character memory and character attributes, both of which align more closely with modern psychological theories.
-We validate the proposed PsyMem using a large scale specially designed role-playing dataset based on real novels, addressing the gap between generic and well-known characters in role-playing tasks.
-Our experiments show that PsyMem-Qwen, trained with our alignment across five dimensions, outperforms all baselines mentioned in this paper, demonstrating strong performance in Character-Independent Capabilities as well.
-Furthermore, we explore the benefits of incorporating Role-based General SFT Data for role-playing models.
-Through this work, we aim to encourage future research on role-playing models that consider psychological frameworks for a more nuanced understanding of character behavior, ultimately advancing the development of more realistic and consistent role-playing LLMs.
+This paper presented PsyMem, a novel framework for LLM-based role-playing that synergizes character memory with fine-grained psychological attributes, aligning with contemporary psychological theories.
+Evaluations conducted on a large-scale novel-based dataset demonstrate PsyMem’s effectiveness in role-playing.
+The resulting model, PsyMem-Qwen, even with only 7B size parameters, outperforms all baselines in character fidelity while also demonstrating strong character-independent capabilities.
+Additionally, we highlighted the advantages of integrating Synthetic Role-playing style General SFT Data.
+We believe this work paves the way for future research into psychologically-informed role-playing models, fostering more nuanced, realistic, and consistent character portrayals and thereby expanding the utility of LLMs in applications such as trustworthy social simulation, enhanced human-computer interaction, and AI counseling.
+Appendix
+Appendix A Source Dataset
+Our dataset was sourced from public English-language novel recommendations on platforms including The StoryGraph (The StoryGraph, 2024), Goodreads (Goodreads, 2024) and BookLikes (BookLikes, 2024).
+An initial pool of 1,710 novels was compiled and subsequently filtered using a Large Language Model (LLM), resulting in a final selection of 539 novels that satisfy the following criteria:
+  • Fluent, reader-friendly novel.
+  • Third-person narration.
+  • Clear formatting and well-structured layout.
+  • Balanced genre coverage to ensure diversity.
+Appendix B Implementation Details
+B.1 Memory Module
+We integrate Memory Module into the PsyMem framework to support contextual memory retrieval during role-playing.
+Specifically, we employed Nano-GraphRAG (Ye, 2024), an open-source lightweight library that implements a simplified version of GraphRAG (Edge et al., 2024).
+Since our test set consists of recently published novels, PsyMem relies on GraphRAG during inference to enrich the contextual content of the dialogue.
+We used a paragraph length of 8192 tokens for the memory text in local mode.
+Other configurations followed the default settings of Nano-GraphRAG.
+Specifically, in each dialogue turn, GraphRAG retrieves relevant entity information mentioned in the prior context to search for memory that aligns with the role the model is playing, and then combines this with character attributes to support inference.
+B.2 Training Details
+We applied the PsyMem method to both Qwen2.5-7B-Instruct and LLaMA3.1-8B-Instruct models.
+During training, we incorporated data augmentation techniques to enhance the models’ ability to understand and generate content in role-playing tasks.
+Specifically, for each training instance, there was a 50% probability of shuffling or masking the role attributes to expand the dataset.
+B.3 Hyperparameters
+All models were trained on four A800 80G GPUs for 3 epochs with a global batch size of 128 and a sequence length of 8,196 (increased to 16,384 during the second training stage).
+LoRA was used with a per-device batch size of 2 and 16 gradient accumulation steps.
+We adopted a cosine learning rate schedule with a warmup ratio of 0.1, and set the learning rate to 1e-4.
+Appendix C Additional Evaluation
+Model Character Fidelity
+Per. Val. SL BD Mem. Avg.
+PsyMem-LLaMA (Rewrite) 80.14 81.47 78.62 74.35 89.80 80.88
+PsyMem-LLaMA (Random) 80.33 81.58 78.37 74.56 90.20 81.00
+PsyMem-LLaMA (Original) 80.47 81.93 78.57 74.47 90.20 81.13
+PsyMem-Qwen(Rewrite) 80.70 81.36 80.20 78.42 91.00 82.34
+PsyMem-Qwen(Random) 80.45 81.63 81.02 78.39 92.00 82.70
+PsyMem-Qwen(Original) 80.40 81.74 80.80 78.48 91.80 82.64
+Table 5: Evaluation of the model’s generalization to input with freer phrasing (Rewrite), shuffled dimension order (Random), and prompts consistent with those used during training (Original).
+C.1 More Flexible Inputs
+To assess PsyMem’s robustness to varied prompt styles, we tested three evaluation settings: the original structured prompt (‘Original’), prompts with randomly shuffled character dimensions (‘Random’), and character profiles rewritten into natural paragraphs by GPT-4o (‘Rewrite’).
+Both PsyMem-LLaMA and PsyMem-Qwen were evaluated across five Character Fidelity dimensions.
+As shown in Table 5, the ‘Random’ variant achieves comparable performance to the ‘Original’ version, suggesting that the model is robust to prompt order perturbations (likely a result of training-time augmentation).
+The ‘Rewrite’ variant results in a slight performance decrease (0.25% for LLaMA, 0.30% for Qwen), potentially due to minor information loss during rephrasing.
+Nonetheless, consistently high Memory scores across all variants demonstrate the robustness and generalization ability of PsyMem under varying input conditions.
+C.2 Evaluation strategies
+To verify the robustness of our evaluation methodology, we compared two scoring strategies: one strategy where each character’s dialogues across different scenes were first concatenated and then assessed collectively, and another strategy where each character’s dialogue, consisting of 15 turns per scene, was evaluated separately before being aggregated.
+We applied both strategies to four top-performing models, using three discriminative evaluators—GPT-4o (Hurst et al., 2024), OpenAI-o3 (OpenAI, 2025), and Gemini-2.5-pro (DeepMind, 2025).
+As shown in Table 6, results across all models and evaluators exhibit strong consistency between the two strategies.
+Notably, for PsyMem-Qwen under GPT-4o evaluation, the average Character Fidelity score under concatenated dialogue input was 82.87%, only 0.23% points higher than the segment-wise evaluation score of 82.64%, indicating that the choice of evaluation strategy does not significantly impact overall conclusions.
+Model GPT-4o o3 Gemini-2.5-pro
+GPT-4o 80.86 79.62 76.92
++ Concatenated Eval. 81.00 79.39 76.70
+Deepseek-R1 82.47 78.40 75.87
++ Concatenated Eval. 82.26 78.66 75.99
+CoSER-70B 82.28 80.23 79.16
++ Concatenated Eval. 81.99 80.40 78.99
+PsyMem-Qwen 82.64 80.74 79.70
++ Concatenated Eval. 82.87 80.43 79.39
+Table 6: Average Character Fidelity scores (Per., Val., SL, BD, Mem.) of advanced role-playing models evaluated by different discriminative models and evaluation approaches.
+“Concatenated Eval.” refers to presenting all dialogues across scenarios simultaneously to the discriminative model for evaluation.
+Metrics Per. Val. SL BD
+Cosine (H $\leftrightarrow$ L) 0.96 0.95 0.96 0.94
+$r_{s}$ (H) 0.74 0.79 0.75 0.77
+$\rho$ (H) 0.72 0.71 0.71 0.70
+Mean (Std) (H) 0.59 $\pm$ 0.47 0.46 $\pm$ 0.42 0.42 $\pm$ 0.39 0.54 $\pm$ 0.41
+$\alpha$ (H) 0.77 0.80 0.74 0.80
+Table 7: Cosine similarity, Pearson correlation coefficient ($\rho$), Spearman’s correlation coefficient ($r_{s}$), and Mean (Std) are used to assess the agreement and score distribution between human raters and GPT-4o across different evaluation dimensions.
+For reference, Krippendorff’s Alpha ($\alpha$) is also reported as a measure of inter-rater reliability among human annotators.
+C.3 Human Evaluation
+To validate our LLM-based evaluation framework, we conducted a large-scale human study involving 40 volunteers who assessed 25 dialogue scenarios (50 characters).
+The process was highly time-intensive, requiring over 10 hours per rater.
+For a robust analysis, we employed several complementary metrics.
+We used Krippendorff’s Alpha ($\alpha$) to measure inter-rater reliability among human annotators, thereby establishing the reliability of our evaluation baseline.
+Concurrently, we assessed the agreement between GPT-4o’s scores and the mean human scores using Cosine Similarity, Pearson’s correlation coefficient ($\rho$), and Spearman’s rank correlation coefficient ($r_{s}$).
+The results, summarized in Table 7, validate our approach.
+We first confirmed a reliable human baseline, with Krippendorff’s Alpha ($\alpha$) values of 0.74–0.80 indicating substantial inter-rater reliability (Krippendorff, 2018; Marzi et al., 2024).
+Against this baseline, GPT-4o’s scores showed strong agreement with human ratings.
+Pearson’s correlation ($\rho$) was consistently strong (0.70–0.72), and Spearman’s rank correlation ($r_{s}$) was similarly high (0.74–0.79) (Fleiss, 1971; Hauke and Kossowski, 2011).
+These results collectively demonstrate that GPT-4o serves as a valid proxy for human evaluation in our task.
+Appendix D Evaluation-stage Input Ablation
+Prompt Set Per. Val. SL BD
+All 80.40 81.74 80.80 78.48
+Per. 78.35 73.33 74.72 73.67
+Val. 74.74 79.66 74.35 74.39
+SL* 68.12 67.86 79.64 69.35
+BD* 69.38 68.77 70.14 77.76
+Base 66.01 65.65 68.05 67.19
+Table 8: Performance of PsyMem-Qwen with different prompt combinations across four evaluation dimensions.
+“All” indicates inclusion of all prompts for the four dimensions; “Base” indicates exclusion of these prompts.
+Bold: Best; Underlined: Second best.
+We further investigated the role of character attributes during inference by varying the input format at evaluation time (as shown Table 8).
+Specifically, we compared three input conditions: (1) a minimal setting with no character dimensions, (2) inputs with only one active dimension, and (3) full inputs containing all four dimensions.
+The model’s performance deteriorated significantly when all dimensions were removed—for example, average scores dropped to as low as 66.01 on the Personality axis and 65.65 on Values.
+In contrast, even a single active dimension led to improved performance on the corresponding evaluation axis; using only the Personality prompt recovered the Personality score to 78.35, and similarly, Values alone yielded 79.66 on its dimension.
+Additionally, we observed that providing Personality and Values information boosted performance on Social & Leadership and Behavioral Decision dimensions (74.72 and 73.67 under Personality-only input), while the reverse was not true.
+This may suggest a hierarchical relationship among these attributes, where latent psychological attributes act as foundational traits that inform and guide performance in the more applied, explicit behavioral patterns.